OpenStudy (anonymous):

solve for d/dx by implicit differentiation: (e^y)(cosx)=1+sin(xy)

OpenStudy (amistre64):

what have you attempted? or what do you not understand about it?

OpenStudy (amistre64):

f g = 1 + h f' g + f g' = h'

OpenStudy (anonymous):

so far i have (e^y)(y')(-sinx)=0+sin(xy) and im not sure if this is right or how to find the derivative of sin(xy)

OpenStudy (amistre64):

(e^y)' (cosx) + (e^y) (cosx)' y' e^y cosx - e^y sinx for the left side

OpenStudy (amistre64):

sin(xy) is a chain rule; h(k) to h'(k) k'

OpenStudy (amistre64):

sin(xy) to cos(xy) * (xy)' cos(xy) * (y+xy')